40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 12017     Entry time: Thu Mar 3 01:25:50 2016     In reply to: 12016     Reply to this: 12019
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap - 2 IR + 1 green beatnotes found 

[ericq, gautam]

Summary of work done tonight:

  • The PDH setup at the Y-end has been restored after I had pulled the whole thing apart some weeks ago to see that nothing was obviously wrong with the uPDH box
  • Adjusted the temperature of the Y-end laser such that a beatnote was obtained - I did this using the IR beat (the end laser temp wasn't updated after the PSL temp was changed recently)
  • The Y green beatnote was found easily, there was no alignment on the PSL table necessary, though there is room to improve this situation (beatnote amplitude was ~ -35dBm though we are used to more like -25dBm)
  • The X green beat remains elusive - I played around with the alignment onto the green beat PD at the PSL table for some time, and the two beams are aligned as far as I can tell given the constrained area available in that area. It may be that I have to clear some optics, do a rigorous near-field/far-field alignment of the two beams and then try again
  • Since we had two strong (-5dBm for Y, -9dBm for X) IR beatnotes, we decided to take the ALS noise spectra for these. So as to not overload the amplifiers, a -10dB attenuator (-6dB) was placed directlty after the Y (X) IR beat PDs, before routing these signals through the usual green beat signal chain. Attached is the measured spectrum. The new values of the temperature sliders at which beatnotes can be found are : 1700 for X and -5990 for Y (spectra taken at these values).

To do:

  • For both ends, find the three temperatures at which we have beatnotes, and choose the middle one
  • PM characterization of AUX X laser - it may be that the excess noise in the X spectrum is due to sub-optimal PDH
  • Align the Y green better at the endtable, also take an OLTF measurement for the Y PDH loop
  • Re-check the alignment onto the green beat PD for the X beat

Remarks:

  • The Lightwave laser controllers differ from the Innolight ones in that it is not possible to directly set the signal to the SLOW control BNC to 0, and have that as the new reference point. Rather, there seems to be some setpoint which is saved as a reference, and the moment any signal is applied to the SLOW control BNC, it adjusts the actual temperature w.r.t. this saved setpoint. I believe it is possible to update this setpoint (it is also possible to update the calibration of the power readout, this is an additional issue at the X end), but since this wasn't critical, I've left it as is for the moment...
  • The ALS nosie spectrum for the Y arm IR beat is surprisingly good!
Attachment 1: IR_beat_20160303_2.pdf  59 kB  | Hide | Hide all
IR_beat_20160303_2.pdf
ELOG V3.1.3-